Opened 5 years ago

Closed 3 years ago

Last modified 3 years ago

#572 closed flight processing (fixed)

RG12/11, flight day 100/2015, Lincolnshire

Reported by: dac Owned by:
Priority: immediate Milestone: 2015 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by dac)

Data location: ~/arsf/arsf_data/2015/flight_data/uk/RG12_11-2015_100_Lincolnshire

Data arrived from ARSF via network transfer on 20/04/2015.

Scientific objective: Generation of high quality land cover maps for input into Land Surface Models (LSMs).

Priority: grade alpha4/medium (8/9)

PI: Mathias Disney

Sensors:

  • Fenix (requested)
  • Leica LIDAR (requested)
  • Leica LiDAR FW (not specified)
  • RCD (requested)

Change History (35)

comment:1 Changed 5 years ago by dac

  • Description modified (diff)

comment:2 Changed 4 years ago by lah

Navigation Processing

Basestation ppp:

Latitude: 53 15 00.93616
Longitude: -0 31 12.34717
Ell. Height: 103.667

comment:3 Changed 4 years ago by lah

Navigation Processing Complete

Can't get a good solution in IPAS TC, so using IPAS Pro. Had to remove ends of the data and set maximum distance for dual frequency to 150 km to get a good solution. Resulting Sol file appears to go back in time in analysis graphs, apl confirms this. Can't remove this by cutting out affected parts of the flightline in IPAS Pro/GrafNav, so have created a new sol file cutting off the data that goes back in time. The hyperspectral files now seem to have processed fine.

The RCD sol file does not appear to have this problem, with the only change being the lever arms.

comment:4 Changed 4 years ago by lah

Navigation Processing

Can get a good solution in IPAS TC using elevation mask of 8 and cutting out most of the extra flight (set start time to 473400.000 and stop to 475508.000). Will use these files as can't resolve IPAS Pro bug.

Last edited 4 years ago by lah (previous) (diff)

comment:5 Changed 4 years ago by lah

Fenix Processing

Found SCTs and checked with OS Open Map -Local:

Flightline FENIX
1 0.99
2 0.97

Waiting for calibration to map files.

comment:6 Changed 4 years ago by dac

LiDAR Processing

Processed with 20150406 calibration file. After reprocessing the boresite flight (2015/96) with the original navigation pitch and roll values were OK.

There are only two lines and they don't overlap making checking pitch and roll difficult. Will proceed with classification.

Last edited 4 years ago by dac (previous) (diff)

comment:7 Changed 4 years ago by dac

LiDAR Processing

Classified LAS files.

comment:8 Changed 4 years ago by dac

LiDAR Processing

Created delivery, still need to generate logsheet and readme.

comment:9 Changed 4 years ago by dac

LiDAR Processing

Logsheet and readme generated - ready for delivery check.

comment:10 Changed 4 years ago by gej

Lidar Delivery Check

Some mis-classified points in Line 01, mainly Trees and Pylons. Line 02 seems fine.
I have renamed .sol file from RG12_11-2015-100.sol to RG12_11-2015_100.sol as proj_tidy was complaining.
Everything else is fine.

Last edited 4 years ago by gej (previous) (diff)

comment:11 Changed 4 years ago by dac

LiDAR

Updated line 01 classification, copied to full waveform and recreated ASCII file.

comment:12 Changed 4 years ago by gej

LiDAR DC
flightline 01 ascii has incorrect format, is ASCII. needs to be ASCII, with CRLF Terminators. Possibly run:
iconv -f ASCII -t ASCII, with CRLF line terminators LDR-RG12_11-2015-100-01.txt
not sure if that's the right command though.

Once that's done, ready for delivery.

comment:13 Changed 4 years ago by dac

LiDAR

Fixed line endings using unix2dos. Marking as ready to delivery.

comment:14 Changed 4 years ago by dac

LiDAR Delivery

Delivered to PI via FTP 2015-10-14

comment:15 Changed 4 years ago by lah

Fenix
Ready for delivery check.

comment:16 Changed 4 years ago by asm

Fenix Delivery Check
Started. Replaced the old version of data quality report with the updated one.

comment:17 Changed 4 years ago by asm

Fenix Delivery Check

-All data is correct.
-Spectra looks fine, good match to 6S simulated spectra.
-Creating zipped files.

DC Finished. Will be ready to delivery when zipped files have been created.

comment:18 Changed 4 years ago by dac

Fenix Delivery Check

  • Also checked spectra - looks fine.
  • There is quite a lot of haze in the images (as noted on the logsheet), have added under quality reports in the readme.

comment:19 Changed 4 years ago by asm

Fenix Delivery Check
Zipped files have been created, ready for delivery

comment:20 Changed 4 years ago by asm

Fenix Delivery
Fenix data delivered via FTP. Notification has been sent.

comment:21 Changed 4 years ago by dac

RCD Processing

Started conversion of raw images using CH39_030_060_27_RGB_20100128 config file for loan RCD.

comment:22 Changed 4 years ago by dac

RCD Processing

Copied processed Tiff images back. Ready for tagging.

comment:23 Changed 4 years ago by gej

RCD processing

Images tagged, making delivery.

comment:24 Changed 4 years ago by gej

RCD Processing

Waiting for data quality report.

comment:25 Changed 4 years ago by dac

RCD Processing

New data quality report created.

comment:26 Changed 4 years ago by gej

RCD Processing

Data quality report added to delivery.
Ready for Delivery Check.

comment:27 Changed 4 years ago by asm

RCD Delivery Check

All checks have been made. One minor problem encountered:
-RCD calibration certificate has a different format naming (gej just updated)

A red line was encountered on all pictures.

Last edited 4 years ago by asm (previous) (diff)

comment:28 Changed 4 years ago by gej

RCD

RCD was originally processed with the wrong calibration file.
I have started processing with correct one now.

comment:29 Changed 4 years ago by gej

RCD

Created new delivery.
Ready for delivery check.

comment:30 Changed 4 years ago by asm

RCD Delivery Check

All checks made, no problems encountered.
A second DC is needed since this is my fist RCD DC.

comment:31 Changed 4 years ago by lah

RCD 2nd Delivery Check
Everything looks ok, so ready to deliver.

comment:32 Changed 4 years ago by lah

PI emailed delivery notification 07/12/15.

comment:33 Changed 3 years ago by dac

Archiving

Started uploading to NEODC.

comment:34 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Data archived at NEODC. Closing ticket.

comment:35 Changed 3 years ago by lah

Camera is available from NEODC so marking that as archived as well.

Note: See TracTickets for help on using tickets.